# DogeRM: Equipping Reward Models with Domain Knowledge through Model Merging

### v1

@@ Acknowledgments
We thank the National Center for High-performance Computing (NCHC) for providing computational and storage resources. We also thank Yen-Ting Lin and Chao-Wei Huang for the valuable discussions.
@@ A.1 Supervised Fine-Tuning (SFT)
We use LlamaFactory Zheng et al. (2024) for supervised fine-tuning (SFT). For fine-tuning the backbone of our LLaMA-2 RM, we use Alpacafarm Dubois et al. (2023) with a learning rate of 1e-5 and a batch size of 128. For code generation, we follow a training procedure similar to Wei et al. (2024). First, we use OSS-Instruct to fine-tune LLaMA-2-7B Touvron et al. (2023) for 2 epochs. Then, we continuously fine-tune the model with Magicoder-Evol-Instruct for 1 epoch. The learning rate for both stages is 1e-5, and the effective batch size is 128.
@@ A.2 Reward Modeling
For reward modeling, we modify the sample code provided by TRL von Werra et al. (2020). We trained the backbone model described in the previous section on UltraFeedback Cui et al. (2023) for 1 epoch, using a learning rate of 1e-5 and a batch size of 32.
Appendix B Prompt Template
@@
You are a helpful, respectful and honest assistant. Always
answer as helpfully as possible, while being safe. Your
answers should not include any harmful, unethical, racist,
sexist, toxic, dangerous, or illegal content.
Please ensure that your responses are socially unbiased
and positive in nature. If a question does not make any
sense, or is not factually coherent, explain why instead
of answering something not correct. If you don’t know the
answer to a question, please don’t share false information.
@@
You are a math problem solver. Please think step by step and
demonstrate your calculation steps. After your reasoning
steps, you should generate the answer by following the for-
mat starting with ’The answer is’
@@
Appendix C Dataset Details
Alpacafarm Dubois et al. (2023)
@@ Appendix C Dataset Details
The Alpacafarm dataset consists of 52k instructions as well as response generated by text-davinci-003 model OpenAI (2023) from the original Alpaca dataset Taori et al. (2023). Alpacafarm splits the datasets into 10k ’sft’ subset for instruction fine-tuning, 10k ’pref’ subset for preference learning, 20k ’unlabeled subset for training such as PPO, and 2k ’val’ subset for validation. We only utilize the 10k ’sft’ subset for fine-tuning the backbone of our reward model.
UltraFeedback Cui et al. (2023)
@@
This dataset consists of 64k prompts from sources including UltraChat Ding et al. (2023), ShareGPT Chiang et al. (2023), Evol-Instruct Xu et al. (2024), TruthfulQA Lin et al. (2022), FalseQA Hu et al. (2023), and FLAN Longpre et al. (2023). The responses are generated by a pool of different LLMs. The preferences are generated by GPT-4 Achiam et al. (2023). In our experiment, we use a cleaned version of UltraFeedback [^1] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned Bartolome et al. (2023), which removes TruthfulQA contamination and uses the average of the preference ratings.
OSS-Instruct & Magicoder Evol-Instruct Wei et al. (2024)
@@
OSS-Instruct consists of 75k synthesized data collected by prompting ChatGPT Achiam et al. (2023) to generate a coding problem and solution based on a seed code snippet from an open-sourced platform. The Magicoder Evol-Instruct dataset, based on the work in Luo et al. (2023), uses an open-source implementation theblackcat102 (2023) that has been further decontaminated, resulting in 110k data points for fine-tuning. Both OSS-Instruct and Magicoder Evol-Instruct are used to fine-tune the Code Model for merging.
RewardBench Lambert et al. (2024)
@@
RewardBench is a benchmark designed to evaluate reward models (RMs). The datasets are categorized into core sets and prior sets. The prior sets consist of testing sets from open-sourced preference dataset such as OpenAI Summarization Stiennon et al. (2020b), Anthropic Helpful split Bai et al. (2022), Anthropic HHH subset Askell et al. (2021), and Stanford Human Preference (SHP) Ethayarajh et al. (2022).
@@
We utilize the core sets for evaluation, which include four categories: chat, chat-hard, reasoning, and safety. The chat category collects data from AlpacaEval Li et al. (2023) and MT Bench Zheng et al. (2023) to assess RMs’ basic ability to discern correct responses in open-ended dialogue. Chat-Hard incorporates data from MT Bench Zheng et al. (2023) with similar ratings and LLMBar Zeng et al. (2024) data designed to challenge LLM-based judges. The reasoning category includes math data selected from PRM800K Lightman et al. (2024), where the prompt is the reference answer and the rejected prompt is a wrong solution generated by GPT-4 Achiam et al. (2023). The coding data utilizes HumanEvalPack Muennighoff et al. (2024), augmenting HumanEval Chen et al. (2021) across six programming languages, with the prompt being the reference solution and the rejected prompt being buggy solutions. Safety category comprises data from XSTest Röttger et al. (2024), Do-Not-Answer Wang et al. (2024b), and an in-development refusals dataset at AI2, aiming to accurately test models’ ability to refuse dangerous content and avoid incorrect refusals triggered by similar words.
Auto-J Eval Li et al. (2024)
@@
Auto-J Eval’s pairwise testing set includes examples from various sources: OpenAI Summarization Stiennon et al. (2020a), WebGPT Nakano et al. (2022), Stanford SHP Ethayarajh et al. (2022), Synthetic GPT-J Havrilla (2023), and PKU-SafeAlignment Ji et al. (2023). GPT-4 Achiam et al. (2023) serves as the annotator. The dataset consists of categories including Summarization, Exam Questions, Code, Creative Writing, Functional Writing, Rewriting, General Communication, and NLP Tasks. We exclude the tied examples and re-group the data into Code, Math (extract from Exam Questions category), and Others, following Yuan et al. 2024.
GSM8K Cobbe et al. (2021)
@@
Appendix D Open-Source Model Details
MetaMath Yu et al. (2024)
We use the LLaMA-2-7B based model fine-tuned by the authors for merging. The MetaMath-7B models are trained on the MetaMathQA dataset, which the authors curated by bootstrapping problems from GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021).
MAmmoTH Yue et al. (2024a)
We merge our RM with the MAmmoTH-7B model, a LLaMA-2-7B based model fine-tuned on the MathInstruct dataset. This dataset combines a diverse range of math problems and hybrid rationales curated by the author.
Mistral-RM Ray2333 (2024)
We utilize a Mistral-based reward model trained on a variety of preference datasets to test our framework’s adaptability. This reward model is initialized from Mistral-7B-Instruct-v0.2. Detailed information about the training setup can be found in the author’s blog [^2] https://www.notion.so/Reward-Modeling-for-RLHF-abe03f9afdac42b9a5bee746844518d0.
MAmmoTH2-Plus Yue et al. (2024b)
To test the adaptability of our framework across different model architectures, we use the MAmmoTH2-7B-Plus and merge it with the Mistral RM. This model is fine-tuned from the MAmmoTH2-7B, which is fine-tuned from Mistral-7B-Instruct-v0.2, on public instruction tuning datasets to further enhance performance.
Appendix E Case Study
We conducted a case study on the predictions of the reward model. We examined the changes in chosen/rejected prompt rewards on the coding subset of RewardBench before and after merging the reward models. We identified that the problems initially predicted incorrectly but corrected after merging can be categorized into math, string operations, and bit operations. Of the four problems correctly predicted before merging but incorrectly predicted after merging, two were related to list/sorting operations, while the others involved mathematical operations such as modular arithmetic and derivatives.
E.1 Passed Case after Merging (Math)
Instruction
Write a Python function ‘max_fill(grid, capacity):
import math‘ to solve the following problem:
You are given a rectangular grid of wells. Each row represents a single well,
and each 1 in a row represents a single unit of water.
Each well has a corresponding bucket that can be used to extract water from it,
and all buckets have the same capacity.
Your task is to use the buckets to empty the wells.
Output the number of times you need to lower the buckets.
@@ E.1 Passed Case after Merging (Math)
E.2 Passed Case after Merging (String)
Instruction
@@ E.2 Passed Case after Merging (String)
Concatenate list of strings into a single string
$>$ concatenate([])
$>$ concatenate([’a’, ’b’, ’c’])
’abc’
@@
E.3 Passed Case after Merging (Bit Operation)
Instruction
@@ E.3 Passed Case after Merging (Bit Operation)
Perform binary XOR on these inputs and return result also as a string.
$>$ string_xor(’010’, ’110’)
’100’
@@
    return ’’.join(xor(x, y) for x, y in
zip(a, b))
E.4 Failed Case after Merging (List/Sorting)
Instruction
@@ E.4 Failed Case after Merging (List/Sorting)
Given an array of non-negative integers, return a copy of the given array after sorting,
you will sort the given array in ascending order if the sum( first index value, last index value) is odd,
or sort it in descending order if the sum( first index value, last index value) is even.
Note:
* don’t change the given array.
@@
E.5 Failed Case after Merging (Math)
Instruction
@@ E.5 Failed Case after Merging (Math)
xs represent coefficients of a polynomial.
xs[0] + xs[1] * x + xs[2] * x^2 + ....
@@
Appendix F Full Results
@@
F.1 Best-of-N
Figure 4 and  5 demonstrate the results of Best-of-N sampling on GSM8K when merging our LLaMA-2 RM with MetaMath-7B Yu et al. (2024) and MAmmoTH-7B Yue et al. (2024a), respectively. DogeRM shows consistent improvement across different models being merged.
Figure 6 shows the result of Best-of-N sampling on MBPP when merging our LLaMA-2 RM with the Code Model. DogeRM improves the reranking accuracy, indicating that our method can be generalized to different tasks.
Finally, Figure 7 shows the results when merging the Mistral RM Ray2333 (2024) with MAmmoTH2-7B-Plus Yue et al. (2024b). DogeRM improves the reranking accuracy at an N=16 setting by 2.88%, indicating that our method can be generalized to different model architectures.
F.2 RewardBench

### v2

@@ Acknowledgments
We thank the reviewers for their insightful comments. This work was financially supported by the National Science and Technology Council (NSTC) in Taiwan, under Grants 111-$2$-E-002-013-MY3 and 112-2223-E002-012-MY5. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources. We are also grateful to Yen-Ting Lin, Wei-Lin Chen, Chao-Wei Huang and Wan-Xuan Zhou from National Taiwan University for their insightful discussions and valuable advice on Figure 1.
@@ A.1 Supervised Fine-Tuning (SFT)
We use LlamaFactory (Zheng et al., 2024) for supervised fine-tuning (SFT). For fine-tuning the backbone of our LLaMA-2 RM, we use Alpacafarm (Dubois et al., 2023) with a learning rate of 1e-5 and a batch size of 128. For code generation, we follow a training procedure similar to Wei et al. (2024). First, we use OSS-Instruct to fine-tune LLaMA-2-7B (Touvron et al., 2023) for 2 epochs. Then, we continuously fine-tune the model with Magicoder-Evol-Instruct for 1 epoch. The learning rate for both stages is 1e-5, and the effective batch size is 128.
@@ A.2 Reward Modeling
For reward modeling, we modify the sample code provided by TRL (von Werra et al., 2020). We trained the backbone model described in the previous section on UltraFeedback (Cui et al., 2024) for 1 epoch, using a learning rate of 1e-5 and a batch size of 32.
For continuous fine-tuning of LLaMA-2 RM on Auto-J Eval (Li et al., 2024) math and code test data, we set the learning rate to 1e-6, the batch size to 8, and the number of epochs to 1.
Appendix B Best-of-N Sampling
We use vLLM (Kwon et al., 2023) to generate responses for reranking. For the GSM8K dataset (Cobbe et al., 2021), we set the temperature to 1.0, top-p to 1.0, and a maximum token length of 512. In the case of MBPP (Austin et al., 2021), we adjust the temperature to 0.1, top-p to 0.95, and maintain the same max length of 512, aligning with the hyperparameters from the bigcode-evaluation-harness [^2] https://github.com/bigcode-project/bigcode-evaluation-harness repository (Zhuo et al., 2024).
Appendix C Prompt Template
@@
You are a helpful, respectful and honest
assistant. Always answer as helpfully as
possible, while being safe. Your answers
should not include any harmful, unethical,
racist, sexist, toxic, dangerous, or illegal
content.
Please ensure that your responses are socially
unbiased and positive in nature. If a question
does not make any sense, or is not factually
coherent, explain why instead of answering
something not correct. If you don’t know the
answer to a question, please don’t share false
information.
@@
You are a math problem solver. Please think
step by step and demonstrate your calculation
steps. After your reasoning steps, you should
generate the answer by following the format
starting with ’The answer is’
@@
Appendix D Dataset Details
Alpacafarm (Dubois et al., 2023)
@@ Appendix D Dataset Details
The Alpacafarm dataset consists of 52k instructions as well as response generated by text-davinci-003 model (OpenAI, 2023) from the original Alpaca dataset (Taori et al., 2023). Alpacafarm splits the datasets into 10k ’sft’ subset for instruction fine-tuning, 10k ’pref’ subset for preference learning, 20k ’unlabeled subset for training such as PPO, and 2k ’val’ subset for validation. We only utilize the 10k ’sft’ subset for fine-tuning the backbone of our reward model.
UltraFeedback (Cui et al., 2024)
@@
This dataset consists of 64k prompts from sources including UltraChat (Ding et al., 2023), ShareGPT (Chiang et al., 2023), Evol-Instruct (Xu et al., 2024), TruthfulQA (Lin et al., 2022), FalseQA (Hu et al., 2023), and FLAN (Longpre et al., 2023). The responses are generated by a pool of different LLMs. The preferences are generated by GPT-4 (Achiam et al., 2023). In our experiment, we use a cleaned version of UltraFeedback [^3] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned (Bartolome et al., 2023), which removes TruthfulQA contamination and uses the average of the preference ratings.
OSS-Instruct & Magicoder Evol-Instruct (Wei et al., 2024)
@@
OSS-Instruct consists of 75k synthesized data collected by prompting ChatGPT (Achiam et al., 2023) to generate a coding problem and solution based on a seed code snippet from an open-sourced platform. The Magicoder Evol-Instruct dataset, based on the work in (Luo et al., 2024), uses an open-source implementation (theblackcat102, 2023) that has been further decontaminated, resulting in 110k data points for fine-tuning. Both OSS-Instruct and Magicoder Evol-Instruct are used to fine-tune the Code Model for merging.
RewardBench (Lambert et al., 2024)
@@
RewardBench is a benchmark designed to evaluate reward models (RMs). The datasets are categorized into core sets and prior sets. The prior sets consist of testing sets from open-sourced preference dataset such as OpenAI Summarization (Stiennon et al., 2020), Anthropic Helpful split (Bai et al., 2022), Anthropic HHH (Askell et al., 2021), and Stanford Human Preference (SHP) (Ethayarajh et al., 2022).
@@
We utilize the core sets for evaluation, which include four categories: chat, chat-hard, reasoning, and safety. The chat category collects data from AlpacaEval (Li et al., 2023) and MT Bench (Zheng et al., 2023) to assess RMs’ basic ability to discern correct responses in open-ended dialogue. Chat-Hard incorporates data from MT Bench (Zheng et al., 2023) with similar ratings and LLMBar (Zeng et al., 2024) data designed to challenge LLM-based judges. The reasoning category includes math data selected from PRM800K (Lightman et al., 2024), where the prompt is the reference answer and the rejected prompt is a wrong solution generated by GPT-4 (Achiam et al., 2023). The coding data utilizes HumanEvalPack (Muennighoff et al., 2024), augmenting HumanEval (Chen et al., 2021) across six programming languages, with the prompt being the reference solution and the rejected prompt being buggy solutions. Safety category comprises data from XSTest (Röttger et al., 2024), Do-Not-Answer (Wang et al., 2024b), and an in-development refusals dataset at AI2, aiming to accurately test models’ ability to refuse dangerous content and avoid incorrect refusals triggered by similar words.
Auto-J Eval (Li et al., 2024)
@@
Auto-J Eval’s pairwise testing set includes examples from various sources: OpenAI Summarization (Stiennon et al., 2020), WebGPT (Nakano et al., 2022), Stanford SHP (Ethayarajh et al., 2022), Synthetic GPT-J (Havrilla, 2023), and PKU-SafeAlignment (Ji et al., 2023). GPT-4 (Achiam et al., 2023) serves as the annotator. The dataset consists of categories including Summarization, Exam Questions, Code, Creative Writing, Functional Writing, Rewriting, General Communication, and NLP Tasks. We exclude the tied examples and re-group the data into Code, Math (extract from Exam Questions category), and Others, following Yuan et al. 2024.
GSM8K (Cobbe et al., 2021)
@@
Appendix E Open-Source Model Details
MetaMath (Yu et al., 2024)
We use the LLaMA-2-7B based model fine-tuned by the authors for merging. The MetaMath-7B models are trained on the MetaMathQA dataset, which the authors curated by bootstrapping problems from GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). According to the original paper, the model did not trained on any data from the testing set of GSM8K and MATH.
MAmmoTH (Yue et al., 2024a)
We merge our RM with the MAmmoTH-7B model, a LLaMA-2-7B based model fine-tuned on the MathInstruct dataset. This dataset combines a diverse range of math problems and hybrid rationales curated by the author. According to the original paper, the model did not trained on any data from the testing set of GSM8K as well as MATH.
Mistral-RM (Ray2333, 2024)
We use a Mistral-based RM, initialized from Mistral-7B-Instruct-v0.2, trained on diverse preference datasets to evaluate our framework’s adaptability. Detailed information about the training setup can be found in the author’s blog. [^4] https://www.notion.so/abe03f9afdac42b9a5bee746844518d0
MAmmoTH2-Plus (Yue et al., 2024b)
To test the adaptability of our framework across different model architectures, we use the MAmmoTH2-7B-Plus and merge it with the Mistral RM. This model is fine-tuned from the MAmmoTH2-7B, which is fine-tuned from Mistral-7B-Instruct-v0.2, on public instruction tuning datasets to further enhance performance. According to the original paper, the model did not trained on any data from the testing set of GSM8K as well as MATH.
Appendix F Code Model Details
To showcase the capabilities of the fine-tuned code model, we assess its performance on two benchmarks: MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021), utilizing BigCodeBench (Zhuo et al., 2024) for evaluation.
Model MBPP Humaneval
LLaMA-2 18.6 12.2
FT on Alpacafarm 21.0 15.9
Code Model 26.2 31.7
Table 4: Performance on two code benchmarks.
Table 4 demonstrates that our fine-tuned Code model outperforms the original Llama2 and the backbone of our reward model in the code benchmarks.
Appendix G Case Study
Instruction (Math)
Write a Python function ‘max_fill(grid, capacity): import math‘ to solve the following problem: You are given a rectangular grid of wells. Each row represents a single well, and each 1 in a row represents a single unit of water. Each well has a corresponding bucket that can be used to extract water from it, and all buckets have the same capacity. Your task is to use the buckets to empty the wells. Output the number of times you need to lower the buckets.
@@ Appendix G Case Study
Instruction (String)
@@
Concatenate list of strings into a single string $>$ concatenate([]) ’’ $>$ concatenate([’a’, ’b’, ’c’]) ’abc’
@@
Instruction (Bit Operation)
@@
Perform binary XOR on these inputs and return result also as a string. $>$ string_xor(’010’, ’110’) ’100’
@@
    return ’’.join(xor(x, y) for x, y in zip(a, b))
Table 5: Passed cases after merging.
Instruction (List/Sorting)
@@
Given an array of non-negative integers, return a copy of the given array after sorting, you will sort the given array in ascending order if the sum( first index value, last index value) is odd, or sort it in descending order if the sum( first index value, last index value) is even.
Note: * don’t change the given array.
@@
Instruction (Math)
@@
xs represent coefficients of a polynomial. xs[0] + xs[1] * x + xs[2] * x^2 + ....
@@
Table 6: Failed cases after merging.
We conducted a case study on the predictions of the reward model. We examined the changes in chosen/rejected prompt rewards on the coding subset of RewardBench before and after merging the reward models. We identified that the problems initially predicted incorrectly but corrected after merging can be categorized into math, string operations, and bit operations. The passed and failed cases are shown in Table 5 and Table 6 respectively. Of the four problems correctly predicted before merging but incorrectly predicted after merging, two were related to list/sorting operations, while the others involved mathematical operations such as modular arithmetic and derivatives.
Appendix H Full Results
@@
H.1 Best-of-N
Figure 4 and  5 demonstrate the results of Best-of-N sampling on GSM8K when merging our LLaMA-2 RM with MetaMath-7B (Yu et al., 2024) and MAmmoTH-7B (Yue et al., 2024a), respectively. DogeRM shows consistent improvement across different models being merged.
Figure 6 shows the result of Best-of-N sampling on MBPP when merging our LLaMA-2 RM with the Code Model. While merging did not lead to a performance decline, the observed improvement is modest. We suspect this is attributable to the low upper bound of reranking performance (represented by the black line), which limits the potential gains from reranking in this task.
Finally, Figure 7 shows the results when merging the Mistral RM (Ray2333, 2024) with MAmmoTH2-7B-Plus (Yue et al., 2024b). DogeRM improves the reranking accuracy at an N=16 setting by 2.88%, indicating that our method can be generalized to different model architectures.
H.2 RewardBench